# Gaussian-Informed Continuum for Physical Property Identification and Simulation

### v1

@@ Abstract
This paper studies the problem of estimating physical properties (system identification) through visual observations. To facilitate geometry-aware guidance in physical property estimation, we introduce a novel hybrid framework that leverages 3D Gaussian representation to not only capture explicit shapes but also enable the simulated continuum to deduce implicit shapes during training. We propose a new dynamic 3D Gaussian framework based on motion factorization to recover the object as 3D Gaussian point sets across different time states. Furthermore, we develop a coarse-to-fine filling strategy to generate the density fields of the object from the Gaussian reconstruction, allowing for the extraction of object continuums along with their surfaces and the integration of Gaussian attributes into these continuums. In addition to the extracted object surfaces, the Gaussian-informed continuum also enables the rendering of object masks during simulations, serving as implicit shape guidance for physical property estimation. Extensive experimental evaluations demonstrate that our pipeline achieves state-of-the-art performance across multiple benchmarks and metrics. Additionally, we illustrate the effectiveness of the proposed method through real-world demonstrations, showcasing its practical utility. Our project page is at https://jukgei.github.io/project/gic.
@@ 1 Introduction
To address these limitations, this paper proposes a novel hybrid solution based on 3D Gaussians [15, 16] and material point method (MPM) [17, 18]. The core strength of this work is that we make use of both explicit shapes from dynamic 3D Gaussian reconstruction and implicit shapes rendered by the Gaussian-informed continuum for physical property estimation.
To generate more precise shapes to reason physical property, we first propose a motion-factorized dynamic 3D Gaussian network to conduct dynamic scene reconstruction. We then extract the continuum from the recovered 3D Gaussians at each frame by leveraging a coarse-to-fine filling strategy to generate the density field of the object progressively. The resulting density fields can be used to sample continuum particles for simulation and extract object surfaces as explicit-shape supervision in physical property estimation. To eliminate the appearance distortion caused by large deformation in PAC-NeRF, we further assign Gaussian attributes to the continuum particles where the opacity and scale attributes are evaluated from the density field. Such Gaussian-informed continuum are able to render object masks during simulation, which can be regarded as an implicit-shape representation to guide the estimation and effectively avoid using inferior rendering results for learning physical properties.
@@
  • We propose a novel hybrid pipeline that takes advantage of the 3D Gaussian representation of the object to both acquire explicit shapes and empower the simulated continuum to infer implicit shapes for physical property estimation.
@@ 2 Related Work
System identification. Understanding the physics laws of the 3D world is beneficial for simulation [31, 32, 6] and manipulation [2, 3, 33]. However, unveiling these properties from visual information is an extremely difficult task due to the ambiguity introduced by incomplete observation and the high degrees of freedom of the scene. Early works [34, 35] study the problem by learning physical properties via interactions. With recent improvements in differentiable physics simulation [17, 18, 36, 37, 38, 39, 40], many methods turn to evaluate the physical properties by comparing the rendering results with 2D ground truth given the prior knowledge about the object geometry. VEO [5] presents a differentiable simulator to learn patterns from 4D reconstruction and force-displacement measurements. Another approach [4] eliminates the dependence of captured forces by proposing an iteration framework between deformation tracking and parameter optimization. While these methods demonstrate promising results, the inferior reconstruction might lead to degraded performance, and the assumption of elastic material restricts the applicability. PAC-NeRF [12] instead proposes a single framework to recover both the unknown geometry and physical properties of deformable objects from multi-view video sequences. However, the inferior geometries and blurry rendered images might have detrimental effects on physical property reasoning. In this work, we adopt MPM as our simulation framework following the approach used in PAC-NeRF due to its ability to simulate a variety of materials [6, 41, 42, 43]. Unlike previous approaches, we utilize dynamic 3D Gaussians to reconstruct explicit 3D geometries and generate simulatable continuum particles. Furthermore, we enhance the particles with Gaussian attributes, facilitating the rendering of implicit 2D shapes, and thereby improving physical parameter estimation.
@@ 4.2 Motion-factorized Dynamic 3D Gaussian Network
In this work, we regard all the Gaussians as isotropic kernels, which has been demonstrated as an efficient way to simplify the model and better reconstruct the scene [6, 48]. We should note that although previous works [29, 49] also perform motion decomposition modeling, our pipeline shows two major differences: 1) instead of modeling each basis with an independent neural network, our module shares a common backbone. Our key observation is that for reconstructing a dynamic object, all points on the object should follow a similar moving tendency, and the final heads of the neural network are sufficient to model the details of different parts of the object; 2) to increase the ability to fit high rank of the dynamic scene [16], we model the motion coefficients as time-variant variables rather than constant Gaussian attributes [29].
@@ 4.4 Geometry-aware Physical Property Estimation
With the Gaussian-informed continuum at initial state $\mathbb{P}_{\tilde{P}}(0)$ and the extracted surfaces $\tilde{S}(t)$ in place, we can employ MPM to perform simulation on the continuum and evaluate the difference in terms of both the explicit and implicit shapes. Concretely, after a rollout by MPM given the current estimation of physical parameters, we obtain a trajectory $P(t)$ with corresponding object surfaces $S(t)$. We thus can render object masks over the trajectory. Then the loss of the current rollout can be computed as:
@@ 5 Experiments
Datasets. To thoroughly assess our proposed method, we employ two sources of data introduced by PAC-NeRF [12] and Spring-Gaus [6]. Concretely, PAC-NeRF contributes two synthetic datasets generated by MLS-MPM framework [18]. Each object in both datasets includes RGB images from 11 distinct viewpoints, with approximately 14 frames per viewpoint. The datasets feature a range of materials, including elastic and plastic objects, granular media, and both Newtonian and non-Newtonian fluids. The first dataset contains 9 objects with different shapes, while the second one consists of 45 cross-shape objects with different initial conditions and ground-truth values of physical properties. The interpretation of the physical parameters is listed in Appendix A.6 and A.7. Spring-Gaus generates a synthetic dataset of elastic objects and collects a real-world dataset containing both static and dynamic scenes. The synthetic data contains 30 frames in each of 10 viewpoints. While the real-world data only contains 3 viewpoints for each object in the dynamic scene, it captures 50-70 images from various viewpoints for the static scene. Moreover, we follow previous works [12, 6] and use the off-the-shelf matting [51] or segmentation [52] techniques to obtain object masks.
Table 1: System Identification Performance on PAC-NeRF Dataset
PAC-NeRF [12] Ours Ground Truth
@@
5.1 Evaluation on PAC-NeRF Synthetic Dataset
@@
Comparison on dynamic reconstruction. In this experiment, we first perform dynamic Gaussian reconstruction on the cross-shaped object dataset using DefGS and our proposed method, respectively. We then employ the same filling strategy on the reconstructed Gaussians at each time state to generate the continuum, which is regarded as the final recovered geometry of the object and used to make comparisons with the oracle shape to compute CD and EMD. Since PAC-NeRF jointly recovers both geometries and physical parameters, we use the final estimated results to generate the trajectory for evaluation. The results, reported in Tab. 2, show that our method outperforms the baselines on both metrics and achieves more precise reconstruction performance on most objects. Specifically, we find that the NeRF representation used by PAC-NeRF usually leads to overly large shape generation. While DefGS performs well on elastic objects, its performance degenerates when modeling objects with large deformations, such as granular media and fluids. Our method can better handle these objects due to the flexibility of trajectory representation.
@@
Comparison on system identification. We evaluate the performance of system identification of the two datasets proposed by PAC-NeRF. For the first dataset, we execute 10 times of our method with different random seeds for each object instance and report the mean value of the estimation results. For the second dataset, we compute the MAE of the parameters for each type of object. To demonstrate the effectiveness of the implicit shape representation, we also conduct experiments on the second dataset by only using masks for supervision on our method, namely “Ours*”. The training details are illustrated in Appendix A.3.
The results, reported in Tab. 1 and Tab. 3, show that the proposed hybrid pipeline can achieve more accurate estimation over a wide range of entries and objects, which demonstrate the effectiveness of the geometry-aware guidance. Fig. 4 visualizes the RGB images rendered by PAC-NeRF and the masks rendered by our method. We can see that when large deformation occurs, the rendered RGB image becomes distorted, while the rendered mask can effectively reduce such effect and get better performance. By leveraging both explicit and implicit shape guidance, our method obtains the best results on most entries. More qualitative results are available in the supplementary video.
5.2 Evaluation on Spring-Gaus Synthetic Dataset
@@ 6 Conclusion and Limitations
This paper proposes a novel solution that leverages the 3D Gaussian representation of objects to acquire explicit shapes while concurrently enabling the simulated continuum to infer implicit shapes to facilitate the estimation of physical properties. A novel motion-factorized dynamic 3D Gaussian framework is proposed to reconstruct precise dynamic scenes. Object surfaces and Gaussian-informed continuum are obtained by utilizing the proposed coarse-to-fine density field generation strategy. Extensive experiments demonstrate the efficacy and applicability of our method.
@@ A.5 Experiment Setting for Spring-Gaus Real-world Dataset
A.6 Physical Properties

### v2

@@ Abstract
This paper studies the problem of estimating physical properties (system identification) through visual observations. To facilitate geometry-aware guidance in physical property estimation, we introduce a novel hybrid framework that leverages 3D Gaussian representation to not only capture explicit shapes but also enable the simulated continuum to render object masks as 2D shape surrogates during training. We propose a new dynamic 3D Gaussian framework based on motion factorization to recover the object as 3D Gaussian point sets across different time states. Furthermore, we develop a coarse-to-fine filling strategy to generate the density fields of the object from the Gaussian reconstruction, allowing for the extraction of object continuums along with their surfaces and the integration of Gaussian attributes into these continuums. In addition to the extracted object surfaces, the Gaussian-informed continuum also enables the rendering of object masks during simulations, serving as 2D-shape guidance for physical property estimation. Extensive experimental evaluations demonstrate that our pipeline achieves state-of-the-art performance across multiple benchmarks and metrics. Additionally, we illustrate the effectiveness of the proposed method through real-world demonstrations, showcasing its practical utility. Our project page is at https://jukgei.github.io/project/gic.
@@ 1 Introduction
To address these limitations, this paper proposes a novel hybrid solution based on 3D Gaussians [15, 16] and material point method (MPM) [17, 18]. The core strength of this work is that we make use of both 3D shapes from dynamic 3D Gaussian reconstruction and 2D shapes rendered by the Gaussian-informed continuum for physical property estimation.
To generate more precise shapes to reason physical property, we first propose a motion-factorized dynamic 3D Gaussian network to conduct dynamic scene reconstruction. We then extract the continuum from the recovered 3D Gaussians at each frame by leveraging a coarse-to-fine filling strategy to generate the density field of the object progressively. The resulting density fields can be used to sample continuum particles for simulation and extract object surfaces as 3D-shape supervision in physical property estimation. To eliminate the appearance distortion caused by large deformation in PAC-NeRF, we further assign Gaussian attributes to the continuum particles where the opacity and scale attributes are evaluated from the density field. Such Gaussian-informed continuum are able to render object masks during simulation, which can be regarded as a 2D-shape representation to guide the estimation and effectively avoid using inferior rendering results for learning physical properties.
@@
  • We propose a novel hybrid pipeline that takes advantage of the 3D Gaussian representation of the object to both acquire 3D shapes and empower the simulated continuum to render 2D shapes for physical property estimation.
@@ 2 Related Work
System identification. Understanding the physics laws of the 3D world is beneficial for simulation [31, 32, 33, 34, 35, 6] and manipulation [2, 3, 36, 37, 38]. However, unveiling these properties from visual information is an extremely difficult task due to the ambiguity introduced by incomplete observation and the high degrees of freedom of the scene. Early works [39, 40] study the problem by learning physical properties via interactions. With recent improvements in differentiable physics simulation [17, 18, 41, 42, 43, 44, 45], many methods turn to evaluate the physical properties by comparing the rendering results with 2D ground truth given the prior knowledge about the object geometry. VEO [5] presents a differentiable simulator to learn patterns from 4D reconstruction and force-displacement measurements. Another approach [4] eliminates the dependence of captured forces by proposing an iteration framework between deformation tracking and parameter optimization. While these methods demonstrate promising results, the inferior reconstruction might lead to degraded performance, and the assumption of elastic material restricts the applicability. PAC-NeRF [12] instead proposes a single framework to recover both the unknown geometry and physical properties of deformable objects from multi-view video sequences. However, the inferior geometries and blurry rendered images might have detrimental effects on physical property reasoning. In this work, we adopt MPM as our simulation framework following the approach used in PAC-NeRF due to its ability to simulate a variety of materials [6, 46, 47, 48]. Unlike previous approaches, we utilize dynamic 3D Gaussians to reconstruct explicit 3D geometries and generate simulatable continuum particles. Furthermore, we enhance the particles with Gaussian attributes, facilitating the rendering of 2D shapes, and thereby improving physical parameter estimation.
@@ 4.2 Motion-factorized Dynamic 3D Gaussian Network
In this work, we regard all the Gaussians as isotropic kernels, which has been demonstrated as an effective way to simplify the model and better reconstruct the scene [6, 53]. We should note that although previous works [29, 54] also perform motion decomposition modeling, our pipeline shows two major differences: 1) instead of modeling each basis with an independent neural network, our module shares a common backbone. Our key observation is that for reconstructing a dynamic object, all points on the object should follow a similar moving tendency, and the final heads of the neural network are sufficient to model the details of different parts of the object; 2) to increase the ability to fit high rank of the dynamic scene [16], we model the motion coefficients as time-variant variables rather than constant Gaussian attributes [29].
@@ 4.4 Geometry-aware Physical Property Estimation
With the Gaussian-informed continuum at initial state $\mathbb{P}_{\tilde{P}}(0)$ and the extracted surfaces $\tilde{S}(t)$ in place, we can employ MPM to perform simulation on the continuum and evaluate the difference in terms of both the 3D and 2D shapes. Concretely, after a rollout by MPM given the current estimation of physical parameters, we obtain a trajectory $P(t)$ with corresponding object surfaces $S(t)$. We thus can render object masks over the trajectory. Then the loss of the current rollout can be computed as:
@@ 5 Experiments
Datasets. To thoroughly assess our proposed method, we employ two sources of data introduced by PAC-NeRF [12] and Spring-Gaus [6]. Concretely, PAC-NeRF contributes two synthetic datasets generated by MLS-MPM framework [18]. Each object in both datasets includes RGB images from 11 distinct viewpoints, with approximately 14 frames per viewpoint. The datasets feature a range of materials, including elastic and plastic objects, granular media, and both Newtonian and non-Newtonian fluids. The first dataset contains 45 cross-shape objects with different initial conditions and ground-truth values of physical properties, while the second one consists of 9 objects with different shapes. The interpretation of the physical parameters is listed in Appendix A.9 and A.10. Spring-Gaus generates a synthetic dataset of elastic objects and collects a real-world dataset containing both static and dynamic scenes. The synthetic data contains 30 frames in each of 10 viewpoints. While the real-world data only contains 3 viewpoints for each object in the dynamic scene, it captures 50-70 images from various viewpoints for the static scene. Moreover, we follow previous works [12, 6] and use the off-the-shelf matting [56] or segmentation [57] techniques to obtain object masks.
5.1 Evaluation on PAC-NeRF Synthetic Dataset
@@
Comparison on dynamic reconstruction. In this experiment, we first perform dynamic Gaussian reconstruction on the cross-shaped object dataset using DefGS and our proposed method, respectively. We then employ the same filling strategy on the reconstructed Gaussians at each time state to generate the continuum, which is regarded as the final recovered geometry of the object and used to make comparisons with the oracle shape to compute CD and EMD. Since PAC-NeRF jointly recovers both geometries and physical parameters, we use the final estimated results to generate the trajectory for evaluation.
The results, reported in Tab. 1, show that our method outperforms the baselines on both metrics and achieves more precise reconstruction performance on most objects. Specifically, we find that the NeRF representation used by PAC-NeRF usually leads to overly large shape generation. While DefGS performs well on elastic objects, its performance degenerates when modeling objects with large deformations, such as granular media and fluids. Our method can better handle these objects due to the flexibility of trajectory representation.
@@
Comparison on system identification. We evaluate the performance of system identification of the two datasets proposed by PAC-NeRF. For the first dataset, we compute the MAE of the parameters for each type of object. To demonstrate the effectiveness of the 2D shape representation, we also conduct experiments on the second dataset by only using masks for supervision on our method, namely “Ours*”. For the second dataset, we execute 10 times of our method with different random seeds for each object instance and report the mean value of the estimation results. The training details are illustrated in Appendix A.3.
The results, reported in Tab. 2 and Tab. 3, show that the proposed hybrid pipeline can achieve more accurate estimation over a wide range of entries and objects, which demonstrate the effectiveness of the geometry-aware guidance. Fig. 4 visualizes the RGB images rendered by PAC-NeRF and the masks rendered by our method. We can see that when large deformation occurs, the rendered RGB image becomes distorted, while the rendered mask can effectively reduce such effect and get better performance. By leveraging both 3D and 2D shape guidance, our method obtains the best results on most entries. More qualitative results are available in the supplementary video.
Table 3: System Identification Performance on PAC-NeRF Dataset
PAC-NeRF [12] Ours Ground Truth
5.2 Evaluation on Spring-Gaus Synthetic Dataset
@@ 6 Conclusion and Limitations
This paper proposes a novel solution that leverages the 3D Gaussian representation of objects to acquire explicit shapes while concurrently enabling the simulated continuum to render 2D shapes to facilitate the estimation of physical properties. A novel motion-factorized dynamic 3D Gaussian framework is proposed to reconstruct precise dynamic scenes. Object surfaces and Gaussian-informed continuum are obtained by utilizing the proposed coarse-to-fine density field generation strategy. Extensive experiments demonstrate the efficacy and applicability of our method.
@@
7 Acknowledgements
This research was supported by the Research Grant Council of the Hong Kong Special Administrative Region under grant number 16203122. We thank Licheng Zhong for providing us with details about real data collection and links for purchasing objects for real-world experiments.
@@ A.1.1 Implementation details
We employ temporal and positional encoding to the time $t$ and position $\mu_{0}$, respectively, to introduce features with various frequencies. Specifically, the encoding module is denoted as $\gamma(x)=\left(\sin(2^{k}\pi x),\cos(2^{k}\pi x)\right)_{k=0}^{L-1}$, where $L=10$ for both $t$ and $\mu_{0}$.
@@ A.5 Experiment Setting for Spring-Gaus Real-world Dataset
A.6 System Identification Result on PAC-NeRF Dataset
Table 7: System Identification Result on PAC-NeRF Dataset
Newtonian (Init. Guess: $\mu=10$, $\kappa=1.0\times 10^{4}$)
Prediction
$\mu=17.43$
$\kappa=218085.89$
$\mu=449.45$
$\kappa=148550.15$
$\mu=160.74$
$\kappa=177748.93$
$\mu=123.07$
$\kappa=223551.76$
$\mu=48.48$
$\kappa=483821.18$
$\mu=40.57$
$\kappa=14668.64$
$\mu=68.24$
$\kappa=336795.04$
$\mu=230.46$
$\kappa=30713.74$
$\mu=563.53$
$\kappa=24459.41$
$\mu=104.92$
$\kappa=139609.16$
Ground Truth
$\mu=19.46$
$\kappa=56075.55$
$\mu=436.62$
$\kappa=152696.25$
$\mu=155.83$
$\kappa=193525.59$
$\mu=121.76$
$\kappa=257356.05$
$\mu=49.09$
$\kappa=518012.47$
$\mu=38.44$
$\kappa=13772.52$
$\mu=64.16$
$\kappa=358237.13$
$\mu=228.71$
$\kappa=11041.06$
$\mu=552.98$
$\kappa=16789.77$
$\mu=106.93$
$\kappa=112569.73$
Non-newtonian (Init. Guess: $\mu=100.0$, $\kappa=1.0\times 10^{5}$, $\tau_{Y}=10$, $\eta=1$)
Prediction
$\mu=13201.66$
$\kappa=218085.89$
$\tau_{Y}=1246.56$
$\eta=4.34$
$\mu=67041.05$
$\kappa=148550.15$
$\tau_{Y}=7679.39$
$\eta=24.21$
$\mu=40092.22$
$\kappa=177748.93$
$\tau_{Y}=4238.86$
$\eta=12.85$
$\mu=35080.68$
$\kappa=223551.76$
$\tau_{Y}=5289.52$
$\eta=5.99$
$\mu=22017.29$
$\kappa=483821.18$
$\tau_{Y}=1037.84$
$\eta=70.84$
$\mu=20142.84$
$\kappa=14668.64$
$\tau_{Y}=3930.36$
$\eta=48.23$
$\mu=26122.95$
$\kappa=336795.04$
$\tau_{Y}=3604.00$
$\eta=5.19$
$\mu=48006.58$
$\kappa=30713.74$
$\tau_{Y}=1441.98$
$\eta=1.22$
$\mu=75068.35$
$\kappa=24459.41$
$\tau_{Y}=1422.56$
$\eta=4.07$
$\mu=32390.85$
$\kappa=139609.16$
$\tau_{Y}=4960.55$
$\eta=6.26$
Ground Truth
$\mu=13209.25$
$\kappa=201566.59$
$\tau_{Y}=1151.42$
$\eta=6.68$
$\mu=65351.08$
$\kappa=171054.03$
$\tau_{Y}=7491.70$
$\eta=26.69$
$\mu=43757.04$
$\kappa=249639.94$
$\tau_{Y}=3964.94$
$\eta=23.27$
$\mu=36027.61$
$\kappa=134751.55$
$\tau_{Y}=5061.12$
$\eta=22.31$
$\mu=19593.71$
$\kappa=121836.33$
$\tau_{Y}=1462.78$
$\eta=38.83$
$\mu=20522.72$
$\kappa=14494.30$
$\tau_{Y}=4153.38$
$\eta=27.24$
$\mu=51549.45$
$\kappa=370317.66$
$\tau_{Y}=3203.67$
$\eta=20.43$
$\mu=121865.90$
$\kappa=32859.59$
$\tau_{Y}=1192.76$
$\eta=10.27$
$\mu=241579.97$
$\kappa=30324.98$
$\tau_{Y}=1251.29$
$\eta=10.62$
$\mu=33764.59$
$\kappa=122896.10$
$\tau_{Y}=4689.16$
$\eta=22.89$
Elasticity (Init. Guess: $E=316227.77$, $\nu=0.25$)
Prediction
$E=23008.38$
$\nu=0.3008$
$E=3013161.44$
$\nu=0.2758$
$E=644436.88$
$\nu=0.2875$
$E=431721.84$
$\nu=0.3024$
$E=106731.23$
$\nu=0.3527$
$E=81726.40$
$\nu=0.1250$
$E=178265.23$
$\nu=0.3291$
$E=1106374.86$
$\nu=0.1731$
$E=4230293.75$
$\nu=0.1583$
$E=339834.14$
$\nu=0.2717$
Ground Truth
$E=21905.91$
$\nu=0.2969$
$E=3844263.25$
$\nu=0.2892$
$E=648705.38$
$\nu=0.3066$
$E=459804.63$
$\nu=0.2476$
$E=103565.68$
$\nu=0.2954$
$E=81467.34$
$\nu=0.1798$
$E=179237.53$
$\nu=0.3456$
$E=1117086.13$
$\nu=0.1829$
$E=3650610.75$
$\nu=0.1684$
$E=340267.00$
$\nu=0.2597$
Plasticine (Init. Guess: $E=10000$, $\nu=0.25$, $\tau_{Y}=1000$)
Prediction
$E=174283.92$
$\nu=0.3008$
$\tau_{Y}=1553.91$
$E=4494502.24$
$\nu=0.2758$
$\tau_{Y}=58973.06$
$E=1607386.28$
$\nu=0.2875$
$\tau_{Y}=17967.93$
$E=1230654.13$
$\nu=0.3024$
$\tau_{Y}=27979.03$
$E=484760.99$
$\nu=0.3527$
$\tau_{Y}=1077.11$
$E=405734.16$
$\nu=0.1250$
$\tau_{Y}=15447.70$
$E=682408.49$
$\nu=0.3291$
$\tau_{Y}=12988.84$
$E=2304631.70$
$\nu=0.1731$
$\tau_{Y}=2079.31$
$E=5635256.43$
$\nu=0.1583$
$\tau_{Y}=2023.68$
$E=1049167.13$
$\nu=0.2717$
$\tau_{Y}=24607.02$
Ground Truth
$E=219037.98$
$\nu=0.2620$
$\tau_{Y}=1521.82$
$E=4974443.00$
$\nu=0.1410$
$\tau_{Y}=64508.23$
$E=3795741.25$
$\nu=0.1842$
$\tau_{Y}=16983.20$
$E=1661460.38$
$\nu=0.1078$
$\tau_{Y}=28502.29$
$E=456356.41$
$\nu=0.2499$
$\tau_{Y}=1070.41$
$E=427719.28$
$\nu=0.1566$
$\tau_{Y}=14766.65$
$E=5106352.50$
$\nu=0.1743$
$\tau_{Y}=12059.57$
$E=1382011.00$
$\nu=0.2312$
$\tau_{Y}=2050.87$
$E=1045986.31$
$\nu=0.2258$
$\tau_{Y}=2131.80$
$E=1144808.88$
$\nu=0.1988$
$\tau_{Y}=24397.29$
Sand (Init. Guess: $\theta_{fric}^{0}=10$)
Prediction $\theta_{fric}^{0}=32.9184$ $\theta_{fric}^{0}=34.4715$ $\theta_{fric}^{0}=29.1305$ $\theta_{fric}^{0}=31.7486$ $\theta_{fric}^{0}=45.2390$
Ground Truth $\theta_{fric}^{0}=30.6577$ $\theta_{fric}^{0}=32.3751$ $\theta_{fric}^{0}=26.8816$ $\theta_{fric}^{0}=29.3458$ $\theta_{fric}^{0}=42.2861$
Each material contains 10 instances (5 for granular material) with various object orientations, initial velocities, and physical parameters.
A.7 Notation of Algorithm 1
Table 8: Notation of Algorithm 1
Operator or symbol Explanation
$\mathbb{P}_{G}(t)$ Gaussian particle set at time $t$
$\tilde{P}(t)$ Sampled continuum particles at time $t$
$\tilde{S}(t)$ Sampled surface particles at time $t$, $\tilde{S}(t)\subset\tilde{P}(t)$
$F(t)$ 3D Density field at time $t$
$Proj$ Operation projecting 3D particles into 2D image indices according to the camera parameters
$Discretize$ Operation mapping particle positions to voxel indices on the density field
$GetPosition$ Operation returning 3D positions of the binary field
A.8 Necessity of 2D mask supervision
To evaluate the necessity of 2D mask supervision, we perform system identification on 45 cross-shaped object instances in the PAC-NeRF dataset by our method but with only object surface supervision. The results are reported in Tab. 9. It is obvious to see that combining both 2D and 3D shapes as supervision can achieve more accurate performance compared to using 3D shapes only. Therefore, we believe that utilizing 2D mask supervision to some extent makes up for the errors introduced by the 3D object surfaces extracted from dynamic 3D Gaussians.
Table 9: System identification with/without mask supervision
Type
Parameters
w/o masks
w/ masks
Newtonian
$\log_{10}(\mu)$
$\log_{10}(\kappa)$
$v$
2.19$\pm$2.90
24.2$\pm$22.2
0.20$\pm$0.08
1.53$\pm$1.31
14.8$\pm$19.2
0.20$\pm$0.07
Non-
Newtonian
$\log_{10}(\mu)$
$\log_{10}(\kappa)$
$\log_{10}(\tau_{Y})$
$\log_{10}(\eta)$
$v$
19.4$\pm$27.7
24.0$\pm$24.8
4.58$\pm$9.11
49.1$\pm$40.5
1.33$\pm$0.54
13.5$\pm$18.2
12.9$\pm$16.8
4.80$\pm$3.92
40.7$\pm$24.6
0.19$\pm$0.09
Elasticity
$\log_{10}(E)$
$\nu$
$v$
2.85$\pm$1.94
3.97$\pm$2.64
0.22$\pm$0.10
2.43$\pm$3.29
2.52$\pm$2.03
0.82$\pm$0.32
Plasticine
$\log_{10}(E)$
$\log_{10}(\tau_{Y})$
$v$
25.6$\pm$27.4
9.04$\pm$2.37
1.16$\pm$0.00
25.6$\pm$29.4
1.67$\pm$1.21
0.22$\pm$0.10
Sand
$\theta_{fric}$
$v$
2.55$\pm$2.03
0.31$\pm$0.18
4.18$\pm$0.52
0.17$\pm$0.05
A.9 Physical Properties